# Foundation Models for the Brain and Body

### v1

@@ 4.3 Regularization of Subject-Dependent Weights
5 Extension to Convolutional Layers
@@
While the formulation in Equation 4 is presented for linear transformations, it can be seamlessly extended to convolutional layers. A convolution operation is itself a linear transformation, where the weight matrix $W$ is replaced by a convolutional kernel $W$. The Subject-Conditioned convolutional layer can thus be expressed as:
@@
6 Data
We conduct our experiment on two Brain Computer Interface datasets from BCI Competition IV [32]. The data are publicly available [^2] http://bnci-horizon-2020.eu/database/data-sets with the MOABB-Package [7, 1]. The BCI Competition IV 2a dataset [4] consists of 22-channel EEG recordings from nine subjects. For each subject, data were collected in two sessions (training and evaluation), with 288 trials per session. The labels are balanced across four motor imagery classes: left hand, right hand, foot, and tongue. The BCI Competition IV 2b dataset [20, 19] consists of 3-channel EEG recordings from nine subjects. For each subject, data were collected across five sessions, three for training and two for evaluation, totaling between 680 and 760 trials. The labels are balanced between the classes: left hand and right hand.
7 Results
We evaluate our proposed method, Subject-Conditioned Layer, by integrating it into two distinct EEG decoding architectures: EEGNeX [6], a compact Convolutional Neural Network (CNN), and the Patched Brain Transformer (PBT) [17], a Vision Transformer (ViT) based architecture [8] that reflect recent trends in the field [18, 16]. These models demonstrate the versatility of Subject-Conditioned Layer across diverse architecture types. All experiments were conducted on the BCI Competition IV-2a and IV-2b datasets. Further details on the experimental setup are in Appendix A.
7.1 Qualitative Analysis
To confirm our hypothesis that our Subject-Conditioned Layer learns to decouple shared from subject-specific information, we visualize the PBT’s latent representations using t-SNE [33], see Figure 1. The visualizations reveal a clear separation of the learned representations. As shown in the figure, embeddings from the shared weights ($W_{\text{general}}$) are unstructured by subject (left), whereas those from the subject-specific weights ($W_{s}$) form distinct, individual-specific clusters (middle). The full model (right) integrates both aspects: it preserves the subject-level clustering while also forming meaningful sub-clusters corresponding to target labels, Figure 1, demonstrating that it retains task-relevant information. This demonstrates that the low-rank corrections $W_{s}$ have effectively isolated the idiosyncratic signatures of each subject. Together, these visualizations provide compelling evidence for our method’s proposed mechanism.
7.2 Numerical Results
To validate our approach, we compare Subject-Conditioned Layer against several modeling strategies, with results summarized in Table 1. All numerical results are reported as the mean and standard deviation of three reproducible runs with random seeds $\{1,2,3\}$.
Next, we demonstrate the benefit of knowledge sharing by comparing our method against the common practice of training fully independent Subject-Specific Models. This baseline trains a separate model from scratch for each individual. In Table 1, we report the mean performance across all subjects (the complete results are provided in Table 4). Subject-Conditioned Layer consistently outperforms the average of those independently trained models. This finding highlights that the shared backbone ($W_{\text{general}}$) enables effective knowledge transfer across subjects, yielding better generalization than isolated training on limited single-subject datasets.
Finally, to validate our specific architectural design, we introduce a crucial parameter control: Subject-Specific LoRA Models. For each subject, a separate low-rank model is trained from scratch. This baseline allows us to test if the performance gain is simply an artifact of having a certain number of trainable parameters per subject. Subject-Conditioned Layer’s superiority in this comparison is a critical finding. It proves that our hybrid architecture is more effective than training isolated, parameter-efficient models. This validates that the structure of our approach is important for its performance.
Table 1: Numerical Results. Performance comparison across model variants: a subject-agnostic model (no personalization), subject-specific models (trained independently per subject), subject-specific LoRA models (entirely low-rank per subject), and Subject-Conditioned Layer.
Method Total Active BCI Comp. IV2a BCI Comp. IV2b
# Params # Params (4 Classes) (2 Classes)
EEGNeX Subject-Agnostic [6] 155 972 55 972 55.00% $\pm$ 0.66% 74.28% $\pm$ 0.39%
EEGNex Subject-Specific LoRA 181 124 10 980 53.60% $\pm$ 2.91% 69.64% $\pm$ 0.27%
EEGNeX Subject-Conditioned Layer 134 884 64 740 64.72% $\pm$ 0.56% 76.48% $\pm$ 0.25%
PBT Subject-Agnostic [17] 1 867 460 867 460 50.82% $\pm$ 0.85% 75.49% $\pm$ 1.10%
PBT Subject-Specific LoRA 1 680 736 135 712 25.81% $\pm$ 0.52% 53.70% $\pm$ 0.80%
PBT Subject-Conditioned Layer 1 480 612 935 588 54.51% $\pm$ 0.30% 76.36% $\pm$ 0.23%
8 Limitations & Outlook
Our work focuses on introducing Subject-Conditioned Layer and empirically demonstrating its effectiveness on the BCI Competition IV 2a and 2b datasets. However, our study has some limitations, which suggest directions for future research.
A limitation is the framework’s dependence on subject-specific data to train the personalized adapters, $W_{s}$. This means that for previously unseen subjects, predictions can only be made using the shared weights, $W_{\text{general}}$. Although we have shown that such personalization enhances performance, further analysis is required to fully assess the model’s generalization capabilities and its viability for practical deployment. Subject-Conditioned Layer is well-suited for a pre-training and fine-tuning paradigm. One could first train the shared weights $W_{\text{general}}$ across a large population. Subsequently, for a new subject, one could freeze the powerful shared model and efficiently fine-tune only a lightweight subject-specific adapter $W_{s}$. This iterative application, coupled with scaling, could lead to a Foundation model for EEG Decoding. To tackle the challenge of unseen subjects, future research could investigate the geometry of the learned adapter space. By comparing the adapters of different subjects (e.g., via their similarity), we could potentially discover subject clusters or archetypes. This could allow us to initialize a new subject adapter from a pre-computed group adapter, enabling faster personalization or even zero-shot prediction, thereby greatly enhancing the practical utility of our method.
Funding Acknowledgements
This project has received funding from the European Regional Development Fund (grants timingMatters and IntelAlgen) under the European Union’s Horizon Europe Research and Innovation Program, from the German Research Foundation DFG within GRK 2297 ‘Mathematical Complexity Reduction’, and from the German Federal Joint Committee (Grant 01VSF23017), as well as German Research Foundation DFG (Grant 537063406), which we gratefully acknowledge.
Appendix A Experiment Design
A.1 EEGNeX

### v2

@@ 4.4 Regularization of Subject-Specific Adapters
5 Decoding Tasks and Datasets
We evaluate our approach on two distinct neural decoding tasks: speech perception from MEG recordings and motor imagery classification from EEG data.
5.1 Speech Perception
We employ the MEG dataset (Gwilliams et al., 2023), a high-quality MEG dataset for natural speech processing. The data set consists of 208-channel MEG recordings from 27 English-speaking participants who listened to four fictional stories from the Masc corpus (Ide et al., 2010). The total recording duration is 56.2h. For each subject, data were collected in two 1-hour sessions, although the second session is missing for five subjects. This yields 4417 training segments and 1363 test segments with a vocabulary of 1810 and 846 unique words, respectively, with 64% word overlap between splits. Following (Défossez et al., 2023b), we use the same data split and preprocessing procedure, and formulate the speech-perception decoding task as a segment retrieval problem: given a short window of MEG signal, the model must identify the corresponding speech segment from a large set of candidate audio segments.
5.2 Motor Imagery
To complement retrieval-based speech decoding with Brain Computer Interface (BCI) classification, we additionally consider two datasets from BCI Competition IV (Tangermann et al., 2012). The BCI Competition IV 2a dataset (Brunner et al., 2008) consists of 22-channel EEG recordings from nine subjects. For each subject, data were collected in two sessions (training and evaluation), with 288 trials per session. The labels are balanced across four motor imagery classes: left hand, right hand, foot, and tongue. The BCI Competition IV 2b dataset (Leeb et al., 2007, 2008) consists of 3-channel EEG recordings from nine subjects. For each subject, data were collected across five sessions, three for training and two for evaluation, totaling between 680 and 760 trials. The labels are balanced between the classes: left-hand and right-hand.
6 Architectures and Training Objectives
To demonstrate the generality of SuLoRA as a drop-in architectural component, we benchmark it across three distinct modeling setups: a convolutional network, a vision transformer-based architecture, and a contrastively trained representation-learning pipeline.
6.1 BrainMagick
BrainMagick (Défossez et al., 2023b) addresses the challenging task of decoding speech from non-invasive MEG and EEG recordings. The architecture consists of two components: a Brain module and a Speech module. The Brain module comprises a spatial attention mechanism that encodes MEG/EEG sensors, a subject-specific linear layer, and a stack of convolutional blocks. The Speech module is a frozen wav2vec 2.0 speech encoder (Baevski et al., 2020). The Brain module is trained using a CLIP (Contrastive Language–Image Pre-Training) loss (Radford et al., 2021) to align latent brain representations with speech representations. A critical design choice in BrainMagick is the use of a subject-specific layer to handle substantial inter-individual variability in MEG/EEG signals. This layer accounts for approximately 2M of the 9.5M (for 27 subjects) trainable model parameters and scales linearly with the number of subjects. We replaced the subject-specific layer of the BrainMagick with SuLoRA.
6.2 EEGNeX
EEGNeX (Chen et al., 2024) represents the state-of-the-art among compact, convolution-only architectures for motor imagery classification, that refines the widely-adopted EEGNet (Lawhern et al., 2018). The network consists of five sequential blocks. Blocks 1 and 2 apply temporal convolutions with batch normalization to learn frequency-selective filter banks, operating analogously to learned Finite Impulse Response filters. Block 3 performs depthwise spatial convolution across electrode channels with max-norm constraints, followed by ELU activation, average pooling, and dropout. Blocks 4 and 5 introduce dilated temporal convolutions to expand the receptive field while maintaining parameter efficiency, with the final block including additional pooling and flattening operations. A linear layer constrained by the maximum norm serves as the classifier. We integrate SuLoRA by replacing standard convolutional layers while preserving depthwise spatial convolution in Block 3.
6.3 Patched Brain Transformer
The Patched Brain Transformer (PBT) (Klein et al., 2025a) adapts the Vision Transformer (Dosovitskiy et al., 2021) architecture for EEG decoding, that reflect recent trends in the field (Kostas et al., 2021; Jiang et al., 2024b). Raw EEG signals are segmented into fixed-duration patches, which are linearly projected into token embeddings. A learned positional embedding encodes both spatial (electrode) and temporal (time window) information, enabling the model to flexibly handle varying sensor configurations and recording durations. The core architecture follows the standard Transformer encoder design. We apply SuLoRA to the linear projections within the Transformer encoder, decomposing the query, key, value, and feed-forward weight matrices into shared and subject-specific components.
7 Results
We evaluate our method on diverse neural signal decoding tasks: speech perception from MEG recordings, and motor imagery classification from EEG, demonstrating the generality of our approach across different brain recording modalities and cognitive paradigms.
Table 1: Speech decoding performance on the (Gwilliams et al., 2023) MEG dataset. We replace the subject-specific layer from BrainMagick (Défossez et al., 2023b) (Baseline) with SuLoRA at varying ranks $r$ and scaling factors $\alpha$. “BL w/o Sub.-Layer” refers to the Baseline where the subject-specific layer is replaced with the same linear layer for all subjects. The parameter count indicates the number of parameters in the subject-specific layer for each respective method. SuLoRA with $r=8$ nearly matches the Baseline performance while reducing parameters by an order of magnitude. SuLoRA with $r=32$ improves upon Baseline results using only half the number of parameters.
Method subject-specific Accuracy [%] $\pm$ Std. [%]
#Params Top 1 Top 5 Top 10
SuLoRA r=4, $\alpha$=4 1,131k 37.56 $\pm$ 0.67 58.72 $\pm$ 0.52 67.06 $\pm$ 0.54
r=4, $\alpha$=8 1,131k 38.85 $\pm$ 0.33 60.17 $\pm$ 0.45 68.52 $\pm$ 0.39
r=4, $\alpha$=12 1,131k 39.04 $\pm$ 0.20 60.36 $\pm$ 0.21 68.77 $\pm$ 0.36
r=8, $\alpha$=8 1,190k 39.82 $\pm$ 0.72 61.03 $\pm$ 0.68 69.39 $\pm$ 0.62
r=8, $\alpha$=16 1,190k 39.61 $\pm$ 0.48 61.02 $\pm$ 0.27 69.40 $\pm$ 0.37
r=8, $\alpha$=24 1,190k 40.32 $\pm$ 0.50 61.5 $\pm$ 0.23 69.79 $\pm$ 0.29
r=16, $\alpha$=16 1,306k 40.50 $\pm$ 0.27 61.77 $\pm$ 0.19 69.97 $\pm$ 0.22
r=16, $\alpha$=32 1,306k 40.50 $\pm$ 0.27 61.86 $\pm$ 0.32 70.09 $\pm$ 0.29
r=16, $\alpha$=48 1,306k 40.68 $\pm$ 0.51 61.94 $\pm$ 0.39 70.20 $\pm$ 0.36
[…]
r=32, $\alpha$=64 1,540k 41.65 $\pm$ 0.76 62.63 $\pm$ 0.84 70.76 $\pm$ 0.76
r=32, $\alpha$=96 1,540k 41.44 $\pm$ 0.32 62.67 $\pm$ 0.31 70.60 $\pm$ 0.18
r=64, $\alpha$=64 1,006k 40.97 $\pm$ 0.44 62.15 $\pm$ 0.49 70.30 $\pm$ 0.49
r=64, $\alpha$=128 1,006k 41.74 $\pm$ 0.49 62.90 $\pm$ 0.43 70.98 $\pm$ 0.45
r=64, $\alpha$=192 1,006k 41.94 $\pm$ 0.92 63.20 $\pm$ 0.65 71.18 $\pm$ 0.61
r=64, $\alpha$=256 1,006k 41.81 $\pm$ 0.21 63.00 $\pm$ 0.11 70.95 $\pm$ 0.21
r=96, $\alpha$=192 1,473k 41.62 $\pm$ 0.13 62.73 $\pm$ 0.17 70.85 $\pm$ 0.16
Baseline 1,968k 40.60 $\pm$ 0.18 61.84 $\pm$ 0.21 70.00 $\pm$ 0.21
BL w/o Sub.-Layer 1,1 73k 19.23 $\pm$ 0.41 36.69 $\pm$ 0.54 45.48 $\pm$ 0.69
Table 2: Zero-Shot. Prediction accuracy (%) under Leave-One-Subject-Out evaluation of EEGNeX (Chen et al., 2024) on BCI Competition IV-2a using SuLoRA, compared to the subject-agnostic baseline. SuLoRA achieves competitive accuracy relative to the baseline when relying solely on the general weights, indicating that general EEG features are effectively captured by these weights (highest values in bold). Transferred Adapters. Prediction accuracy (%) under Leave-One-Subject-Out evaluation when transferring an adapter trained on one training subject to predict the held-out subject. Adapter $i$ corresponds to subject $i$; therefore, diagonal entries are omitted. Transferred adapters outperform the baseline, enabling a form of few-shot learning by either selecting an appropriate adapter or using it as an initialization for LoRA fine-tuning. (Values shown in bold when higher than the Zero-Shot results for SuLoRA and baseline.)
Method Leave Out Subject
Accuracy [%] $\pm$ Std. [%]
Zero-Shot
S1 S2 S3 S4 S5 S6 S7 S8 S9
[…]
EEGNeX SuLoRA 37.04 $\pm$ 4.02 25.23 $\pm$ 0.16 56.37 $\pm$ 4.52 32.76 $\pm$ 0.16 25.12 $\pm$ 0.16 28.13 $\pm$ 1.50 28.93 $\pm$ 0.59 58.22 $\pm$ 1.99 45.14 $\pm$ 2.04
Adapter Transferred Adapters
S1 S2 S3 S4 S5 S6 S7 S8 S9
1 n/a 28.01 $\pm$ 0.82 67.94 $\pm$ 3.23 34.14 $\pm$ 1.61 24.65 $\pm$ 0.28 29.28 $\pm$ 1.89 34.49 $\pm$ 1.93 46.41 $\pm$ 4.03 34.95 $\pm$ 0.87
2 45.83 $\pm$ 3.71 n/a 28.59 $\pm$ 1.66 31.71 $\pm$ 2.05 28.70 $\pm$ 2.90 26.39 $\pm$ 0.57 16.43 $\pm$ 2.36 31.25 $\pm$ 3.97 24.19 $\pm$ 1.66
3 61.81 $\pm$ 3.00 28.12 $\pm$ 0.85 n/a 27.43 $\pm$ 0.98 26.97 $\pm$ 1.56 31.14 $\pm$ 1.45 29.98 $\pm$ 1.56 48.84 $\pm$ 5.42 39.01 $\pm$ 3.70
4 47.34 $\pm$ 5.57 22.92 $\pm$ 1.02 33.91 $\pm$ 5.70 n/a 26.27 $\pm$ 1.31 31.60 $\pm$ 2.32 22.80 $\pm$ 2.54 54.86 $\pm$ 2.60 34.26 $\pm$ 2.01
[…]
6 36.46 $\pm$ 0.29 25.46 $\pm$ 2.27 35.88 $\pm$ 6.29 37.96 $\pm$ 0.16 25.00 $\pm$ 0.29 n/a 30.21 $\pm$ 1.98 45.95 $\pm$ 2.87 43.52 $\pm$ 4.61
7 52.43 $\pm$ 3.00 24.65 $\pm$ 2.14 45.95 $\pm$ 6.12 26.62 $\pm$ 0.87 25.35 $\pm$ 0.00 26.62 $\pm$ 1.61 n/a 32.18 $\pm$ 0.91 32.18 $\pm$ 3.99
8 54.51 $\pm$ 2.88 29.05 $\pm$ 2.87 59.84 $\pm$ 5.67 43.06 $\pm$ 2.42 24.31 $\pm$ 0.57 28.94 $\pm$ 2.01 28.12 $\pm$ 3.07 n/a 45.14 $\pm$ 2.47
9 29.05 $\pm$ 0.59 25.35 $\pm$ 0.28 31.25 $\pm$ 1.99 26.27 $\pm$ 1.56 25.00 $\pm$ 0.00 28.36 $\pm$ 1.61 23.26 $\pm$ 2.47 41.09 $\pm$ 4.13 n/a
7.1 Application to Speech Decoding
To demonstrate that SuLoRA can serve as an efficient drop-in replacement, we reproduced the speech decoding results from  (Défossez et al., 2023b) on the MEG dataset (Gwilliams et al., 2023) (see Sec. 5.1). We used the publicly available BrainMagick GitHub repository (Défossez et al., 2023a) with identical hyperparameters and random seeds {2036, 2037, 2038}. Unlike the original work, which used gradient accumulation across two (Nvidia V100) GPUs, we trained the model on a single (Nvidia H100) GPU. Since contrastive losses draw negative samples only within-GPU batches, our single-GPU configuration provides access to the full batch of negatives.
In Table 1, we compare SuLoRA with the BrainMagick model (Baseline), that achieves $40.60\%$ a top-1 accuracy (top-10: $70.00\%$) using 1,968k parameters in the subject-specific layer. Replacing the subject-specific layer with the same linear layer for all subjects, causes performance to collapse to top-1 accuracy of $19.23\%$ (top-10: $45.48\%$), underscoring the importance of modeling inter-subject variability in neural decoding tasks.
SuLoRA provides a favorable accuracy-efficiency trade-off across the parameter spectrum. For $r=8$ and $\alpha=24$, it closely matched the baseline accuracy of $40.32\%$ top-1 (top-10: $69.79\%$) with only 190k parameters, a reduction of an order of magnitude compared to the baseline. Peak performance is achieved at $r=64$ with $\alpha=192$, reaching $41.94\%$ top-1, $63.20\%$ top-5, and $71.18\%$ top-10 accuracy. Notably, this configuration still uses roughly half the parameters of the baseline (1,006k vs. 1,968k). Further increasing rank does not improve performance. The scaling factor $\alpha$ plays a crucial role in performance. For a fixed rank, we observe consistent improvements when increasing $\alpha$ up to $\alpha=3r$, after which performance plateaus or slightly decreases.
7.2 Leave-One-Subject-Out
Figure 2: Latent Representations Colored by Subject ID vs. Target Labels. The top row compares t-SNE projections of representations derived from (Left) the shared weights ($W_{\text{general}}$), (Middle) the subject-specific weights ($W_{s}$), and (Right) the fused representations from the complete SuLoRA. In the top row, all points are colored by subject ID. The bottom row presents the same data points, but colored according to their ground-truth target labels. This comparison clearly shows that the sub-clusters formed by our full model (top right) correspond directly to the target classes (bottom right), demonstrating that the model learns discriminative, task-relevant features within each subject’s representation space. All representations are obtained from the classification token predicted by the Patched Brain Transformer, with a linear classification head used for visualization clarity.
Table 2 reports prediction accuracy under a Leave-One-Subject-Out (LOSO) evaluation on the BCI Competition IV-2a dataset using EEGNeX. In each fold, one subject is held out for testing while the model is trained on the remaining eight subjects. We compare a subject-agnostic baseline, SuLoRA in a zero-shot setting, and transferred adapters trained on individual source subjects.
In the zero-shot regime, predictions using SuLoRA rely solely on the shared, subject-independent weights. SuLoRA achieves accuracy comparable to the subject-agnostic baseline across most held-out subjects. While performance varies across individuals, SuLoRA matches or exceeds the baseline for some subjects (e.g., Subjects 3 and 9), indicating that the shared representations capture meaningful and transferable EEG structure even without subject-specific adaptation.
We further assess the transferability of subject-specific adapters by applying adapters trained on a source subject to unseen target subjects. Transferred adapters consistently outperform both zero-shot SuLoRA and the subject-agnostic baseline for many subject pairs. Notably, certain adapters yield substantial improvements for specific held-out subjects (e.g., Adapter 1 for Subject 3 and Adapter 8 for Subject 4), highlighting the presence of cross-subject structure that can be exploited without retraining the full model.
Overall, these results demonstrate that SuLoRA supports zero-shot generalization for cross-subject adaptation, while the results of transferred adapters indicate a few-shot personalization without retraining or initialization for LoRA fine-tuning.
Table 3: Numerical Results. Performance comparison across model variants: a subject-agnostic model (no personalization), subject-specific models (trained independently per subject), LoRA (independent low-rank model per subject), and SuLoRA.
Method Total Active BCI Comp. IV2a BCI Comp. IV2b
# Params # Params (4 Classes) (2 Classes)
EEGNeX Subject-Agnostic (Chen et al., 2024) 155 972 55 972 55.00% $\pm$ 0.66% 74.28% $\pm$ 0.39%
EEGNex LoRA 181 124 10 980 53.60% $\pm$ 2.91% 69.64% $\pm$ 0.27%
EEGNeX SuLoRA 134 884 64 740 64.72% $\pm$ 0.56% 76.48% $\pm$ 0.25%
PBT Subject-Agnostic (Klein et al., 2025a) 1 867 460 867 460 50.82% $\pm$ 0.85% 75.49% $\pm$ 1.10%
PBT LoRA 1 680 736 135 712 25.81% $\pm$ 0.52% 53.70% $\pm$ 0.80%
PBT SuLoRA 1 480 612 935 588 54.51% $\pm$ 0.30% 76.36% $\pm$ 0.23%
7.3 Representation Structure and Subject Conditioning
Figure 2 presents t-SNE visualizations (van der Maaten and Hinton, 2008) of the learned representations to analyze the effect of subject conditioning on the latent space. We compare embeddings derived from the general weights $W_{\text{general}}$, the subject-specific weights $W_{s}$, and the fused representations produced by the complete SuLoRA layer. All embeddings are extracted from the classification token predicted by the Patched Brain Transformer, using a linear classification head for visualization. The top row of Figure 2 colors embeddings by subject identity, while the bottom row colors the same points by ground-truth target labels. Representations obtained from the general weights exhibit substantial overlap across subjects, with limited separation by either subject or class. Subject-specific representations show clearer grouping by subject identity, but this structure does not consistently align with the target labels. In contrast, the fused representations produced by the full SuLoRA display well-defined sub-clusters within each subject’s embedding space. These sub-clusters correspond closely to the ground-truth target labels, as evidenced by the alignment between the top-right and bottom-right panels of Figure 2. This indicates that the proposed model learns task-relevant, discriminative features while preserving subject-dependent structure in the representation space. Overall, these visualizations provide qualitative evidence that SuLoRA improves the organization of the latent space by reducing inter-subject variability while enhancing class separability within each subject.
7.4 Application to Motor Imagery EEG Decoding
We evaluate the SuLoRA on motor imagery EEG decoding by integrating it into two distinct architectures: EEGNeX (see Sec. 6.2) and the Patched Brain Transformer (see Sec. 6.3). The following experiment is conducted on the BCI Competition IV-2a and IV-2b datasets (see Sec. 5.2). All numerical results are reported as the mean and standard deviation of three reproducible runs with random seeds $\{1,2,3\}$. Further details on the experimental setup are provided in Appendix B.1.
Next, we demonstrate the benefit of knowledge sharing by comparing our method against the common practice of training fully independent Subject-Specific Models. This baseline trains a separate model from scratch for each individual. In Table 3, we report the mean performance across all subjects (the complete results are provided in Table 7). SuLoRA consistently outperforms the average of those independently trained models. This finding highlights that the shared backbone ($W_{\text{general}}$) enables effective knowledge transfer across subjects, yielding better generalization than isolated training on limited single-subject datasets.
Finally, to validate our specific architectural design, we introduce a crucial parameter control: LoRA EEGNex. For each subject, a separate low-rank model is trained from scratch. This baseline allows us to test if the performance gain is simply an artifact of having a certain number of trainable parameters per subject. SuLoRA’s superiority in this comparison is a critical finding. It proves that our hybrid architecture is more effective than training isolated, parameter-efficient models. This validates that the structure of our approach is important for its performance.
8 Conclusion
We presented the Subject-Specific Low-Rank Adapter, a method for multi-subject neural signal decoding that explicitly separates shared and subject-specific sources of variation. The method integrates into existing architectures with minimal modification, requiring only replacement of standard linear or convolutional layers.
SuLoRA addresses a fundamental obstacle to scaling foundation models for brain decoding: inter-subject distribution shift. The method decomposes network weights into a general component that is learned across all subjects and lightweight, low-rank subject-specific adapters. Our experiments across MEG speech perception and EEG motor imagery tasks demonstrate that SuLoRA consistently outperforms both subject-agnostic baselines and independently trained subject-specific models. In speech decoding, SuLoRA achieves superior accuracy while reducing subject-specific parameters by half compared to previous studies. In the motor imagery classification task, it improves performance for both CNN and transformer architectures across BCI datasets.
The visualization of learned representations confirms that SuLoRA organizes the latent space by reducing inter-subject variability while enhancing class separability within each subject.
More broadly, the design underlying SuLoRA may extend to other multi-source learning problems characterized by systematic distribution shifts across data sources.
9 Limitations
By design, SuLoRA requires subject identity at both training and inference time. This prevents direct application to entirely unseen subjects without adaptation. Although our transferred adapter experiments suggest that cross-subject structure exists, leveraging this for true zero-shot generalization to new subjects remains future work.
The selection of the appropriate rank $r$ and the scaling factor $\alpha$ requires parameter tuning, and optimal values can vary between tasks and architectures. Our experiments suggest $\alpha\approx 3r$ works well, but an exact selection criterion remains an open question.
Our evaluation focuses on datasets with 9–27 subjects. Although the results are consistent across different modalities and architectures, validating SuLoRA on a larger scale, such as with hundreds of subjects, would further strengthen the motivation for the foundation model.
Finally, the total parameter count increases with the number of subjects, although modestly due to the low-rank constraint. For applications involving a substantial number of subjects, it may be essential to employ strategies like sharing adapters among similar subjects or utilizing hierarchical adapter structures
Impact Statement
This work aims to advance the fields of machine learning and neuroscience, in particular by improving the generalization of models across individuals in neural signal decoding. The proposed methods are intended to increase robustness and sample efficiency in learning from human brain data, which benefit research in brain–computer interfaces and cognitive neuroscience. The work does not introduce new data collections, and we do not anticipate negative ethical or societal consequences beyond those commonly associated with machine learning methods applied to brain signals.
Acknowledgments
This project has received funding from the German Federal Joint Committee (Grant 01VSF23017), from the European Regional Development Fund (grants timing Matters and IntelAlgen) under the European Union’s Horizon Europe Research and Innovation Program, and from the German Research Foundation DFG (grants GRK 2297 and 537063406), which we gratefully acknowledge.
The work of Steffen Schotthöfer is sponsored by the Applied Mathematics Progrm at the Office of Advanced Scientific Computing Research, U.S. Department of Energy, and performed at the Oak Ridge National Laboratory, which is managed by UT-Battelle, LLC under Contract No. DE-AC05-00OR22725 with the U.S. Department of Energy. The United States Government retains and the publisher, by accepting the article for publication, acknowledges that the United States Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this manuscript, or allow others to do so, for United States Government purposes. The Department of Energy will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan (http://energy.gov/downloads/doe-public-access-plan). This manuscript has been authored by UT-Battelle, LLC under Contract No. DE-AC05-00OR22725 with the U.S. Department of Energy. The United States Government retains and the publisher, by accepting the article for publication, acknowledges that the United States Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this manuscript, or allow others to do so, for United States Government purposes. The Department of Energy will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan(http://energy.gov/downloads/doe-public-access-plan).
Appendix A Extension to Convolutional Layers
@@
While the formulation in Equation 4 is presented for linear transformations, it can be seamlessly extended to convolutional layers. A convolution operation is itself a linear transformation, where the weight matrix $W$ is replaced by a convolutional kernel $W$. The SuLoRA layer can thus be expressed as:
@@
Appendix B Numerical Experiments
B.1 Experiment Design
EEGNeX